OpenStudy (kingstone):

The formula A = 118e^(0.024t) models the population of a particular city, in thousands, t years after 1998. When will the population of the city reach 140 thousand? A. 2008 B. 2005 C. 2006 D. 2007 Any kind of help is appreciated!

OpenStudy (kingstone):

I think the answer is C. 2006

OpenStudy (phi):

how did you get 2006?

OpenStudy (kingstone):

118e^(0.024t) switch t with 8 because 1998+8=2006 142.977 thousand. In the year 2005 you get an answer under 140, so that leaves 2006 the closest year.

OpenStudy (phi):

I would do 140= 118 e^(0.024t) divide both sides by 118 140/118 = e^(0.024 t) take the natural log ( abbreviated ln) of both sides ln(140/118)= 0.024 t divide both sides by 0.024 ln(140/118)/0.024 = t t = 7.12 years So the population reaches 140K in 1998+7.12 or 2005.12 so in the early part of 2005

OpenStudy (kingstone):

is 2005.12 considered part of the year 2005 or 2006? remember the answer can only be 2005 or 2006, not in between.

OpenStudy (kingstone):

2005.12 is 2006.. so C is the answer in my opinion.